Thursday 14 September 2017

DNS Blip

There was a brief (approx 30 seconds) DNS outage around 12:00 BST (GMT+1) today. An update to DNS access controls went wrong. This was noticed immediately and fixed. All services are now functioning normally. Please contact support if you have any issues.

Friday 28 July 2017

MSO: Several Services Affected [Resolved]

From 16:23 until 17:03 there was an outage which affected several of our datacentre services. VOIP, some web and email services are know to have been affected. Unfortunately this also took down our phone systems, resulting in calls being diverted to staff mobiles. Cause: one of our virtualisation head units rebooted unexpectedly, taking down all related services. We are still investigating the cause. All services are now back online. If you are still experiencing any issues, please contact or call 0131 510 0340.

Friday 28 April 2017

Major Service Outage: various services affected

We're having some network issues with out upstream that is affecting some services. Some web and email services are affected, as well as a small number of customer VMs. Will update this space as soon as we have more information.